# Ab Initio theory of Electron-phonon-coupling-induced Giant Magnetic Moments of Chiral Phonons in Magnetic Materials

### v2

@@ Abstract
Through first-principles calculations guided by the inertially decoupled model, we identified five hexagonal $P6/mmm$ magnets, specifically $\mathrm{EuSi_{2}}$, $\mathrm{GdSi_{2}}$ [57], $\mathrm{EuGa_{2}}$, $\mathrm{GdGa_{2}}$ [58], and $\mathrm{CrB_{2}}$ [59, 60], exhibiting intrinsic phonon Chern states via EPC-enhanced Zeeman splittings ($\Delta_{\text{z}}=2|E_{\text{Z}}|$). As summarized in Table I, they all have large observable phonon gaps at the $\mathrm{K}$ point defined as $\Delta_{\text{ph}}=\Delta_{\text{z}}-\gamma$, where $\gamma$ is phonon linewidth resulting from the EPC [61, 62]. Note that the Zeeman splitting should be larger than the phonon linewidth to produce an observable phonon gap.
@@
Here, we take $\mathrm{EuSi_{2}}$ as a representative example (see SM [44] for details on all candidate materials). The lattice structures of $\mathrm{EuSi_{2}}$ are illustrated in Figs. 3(a, e). Nonmagnetic $\mathrm{Si}$ atoms form a honeycomb-lattice layer, stacked alternately with a magnetic $\mathrm{Eu}$ triangular-lattice layer, along the [001] direction. The atomic masses of $\mathrm{Si}$ and $\mathrm{Eu}$ are approximately $28\,u$ and $152\,u$ ($u$ is atomic mass unit), respectively, with each $\mathrm{Eu}$ atom carrying an out-of-plane magnetic moment of $7\,\mu_{B}$. As a result, $\mathrm{EuSi_{2}}$ can be aptly described by our proposed inertially decoupled model.
@@
Our first-principles calculations reveal that the EPC induces a pronounced Zeeman splitting of chiral phonons in the high-frequency phonon dispersion of ferromagnetic $\mathrm{EuSi_{2}}$ [see Figs. 3(c, d)], opening a sizable full gap ($\approx 0.35$ THz) throughout the Brillouin zone. Given that the Chern number remains $|C|=1$ for all 2D $k_{z}$ slices, $\mathrm{EuSi_{2}}$ realizes a 3D layered phonon Chern state. This is further supported by the presence of boundary modes in the phonon dispersions of $\mathrm{EuSi_{2}}$ under open-boundary conditions, as illustrated in Fig. 3(e). To achieve a 2D Chern state, we construct a sandwiched slab featuring a $\mathrm{Si}$ sublayer between two $\mathrm{Eu}$ sublayers, as shown by the black rectangle in Fig. 3(b). By comparing Figs. 2(f) and 2(g), we can see that the Zeeman splitting generates a substantial topologically nontrivial phonon gap ($\approx 0.30$ THz) with a Chern number of $|C|=1$ in the phonon dispersion of this sandwiched slab, and the accompanying single pair of edge modes traversing the phonon gap, as shown in Fig. 3(h), provide definitive evidence for intrinsic phonon Chern states.
@@
Table 1: The direct phonon gaps at K points for the five candidate materials, where $\gamma$ and $\Delta_{\text{z}}$ represent the phonon linewidth and the phonon Zeeman splitting, respectively. The observable phonon gap $\Delta_{\text{ph}}$ is defined as $\Delta_{\text{z}}-\gamma$.
@@
Discussion and conclusion. The intrinsic phonon Chern state, which hosts topologically protected edge phonon currents, holds significant potential for applications due to its operation without requiring extreme external strong magnetic fields, Coriolis forces, or stress gradients. One potential advantage may be its ability to detect neutral particles by converting them into phonon edge currents. Neutral particles, especially dark matter particles, rarely participate in electromagnetic interactions, making their detection extremely challenging. However, if the energy transferred from the neutral particle to nucleus during the scattering or absorption falls within that of the first “conduction” band above the topological gap in the phonon dispersion, this process may potentially excite the corresponding phonon modes [63, 64], driving the system out of equilibrium. As the system relaxes back to equilibrium, these high-frequency modes gradually shift to lower frequencies and necessarily cross the phonon Chern gap via edge states, producing a edge phonon current immune to scattering[Fig. 4(a)]. This robust edge current is significantly more measurable than conventional phonon excitations dispersed throughout the material. However, the operating temperature of detectors must be sufficiently low to prevent thermal excitation from overcoming the phonon gap, which would otherwise disrupt these edge currents. Thus, the gap determines the maximum operating temperature for such a detector. Figure 4(b) lists gaps and transfered energy ranges for proposed candidate materials (see SM [44] for details), with transfered energy spanning approximately $20\sim 110\,\mathrm{meV}$. This energy range corresponds to low-energy dark matter, therefore our work provides a potential detection mechanism for this recently emphasized category of dark matter candidates [63, 64].
@@
Figure 4: Detectors of neutral particles based on the intrinsic phonon Chern state. (a) Schematic illustrations of the detection process. Neutral particles excite phonon modes in high-frequency “conduction” bands. These modes then gradually relax to lower-frequency regimes via chiral edge phonon channels. (b) Transfered energy ranges and phonon Chern gaps for all the layered candidate materials.
@@
In summary, we have established an ab initio theory enabling systematic predictions of EPC-driven phonon magnetic properties and TRS-breaking phonon dispersions across for insulator, semimetal and metal systems. Based on our theory, we build the inertially decoupled model to seek intrinsic phonon Chern states, and propose candidate materials exhibiting EPC-driven giant topological phonon gaps exceeding 0.3 THz. Furthermore, we demonstrate the potential capability of these topological chiral phonon states to detect neutral particles through robust phonon edge currents. These advances could open broad pathways for investigating phonon magnetic effects in quantum materials.

### v3

@@ Abstract
Through first-principles calculations guided by the inertially decoupled model, we identified five hexagonal ($P6/mmm$) magnets exhibiting intrinsic phonon Chern states, namely, $\mathrm{EuSi_{2}}$, $\mathrm{GdSi_{2}}$ [59], $\mathrm{EuGa_{2}}$, $\mathrm{GdGa_{2}}$ [60], and $\mathrm{CrB_{2}}$ [61, 62]. Remarkably, they all have large observable topological phonon gaps at the $\mathrm{K}$ point defined as $\Delta_{\text{ph}}=\Delta_{\text{z}}-\gamma$, where $\Delta_{\text{z}}=2|E_{\text{Z}}|$ is the phonon Zeeman splitting and $\gamma$ is the phonon linewidth resulting from EPC [63, 64], as summarized in Table I. Note that the Zeeman splitting should be larger than the phonon linewidth to produce an observable phonon gap.
@@
Here, we take $\mathrm{EuSi_{2}}$ as a representative example (see SM [44] for details on all candidate materials). The lattice structures of $\mathrm{EuSi_{2}}$ are illustrated in Figs. 3(a, b). Nonmagnetic $\mathrm{Si}$ atoms form a honeycomb-lattice layer, stacked alternately with a ferromagnetic $\mathrm{Eu}$ triangular-lattice layer, along the [001] direction. The atomic masses of $\mathrm{Si}$ and $\mathrm{Eu}$ are approximately $28\,u$ and $152\,u$ ($u$ is atomic mass unit), respectively, with each $\mathrm{Eu}$ atom carrying an out-of-plane magnetic moment of $7\,\mu_{B}$. As a result, $\mathrm{EuSi_{2}}$ can be aptly described by our proposed inertially decoupled model.
@@
Our first-principles calculations reveal that the EPC induces a pronounced Zeeman splitting of chiral phonons in the high-frequency phonon dispersion of ferromagnetic $\mathrm{EuSi_{2}}$ [see Figs. 3(c, d)], opening a sizable full topological gap ($\approx 0.35$ THz) throughout the Brillouin zone. Given that the Chern number remains $|C|=1$ for all 2D $k_{z}$ slices, $\mathrm{EuSi_{2}}$ realizes a 3D layered phonon Chern state. This is further supported by the presence of boundary modes in the phonon dispersions of $\mathrm{EuSi_{2}}$ under open-boundary conditions, as illustrated in Fig. 3(e). To achieve a 2D Chern state, we construct a dynamically-stable sandwiched slab featuring a $\mathrm{Si}$ sublayer between two $\mathrm{Eu}$ sublayers, as shown by the black rectangle in Fig. 3(b). By comparing Figs. 2(f) and 2(g), we can see that the phonon Zeeman splitting generates a substantial topologically nontrivial phonon gap ($\approx 0.30$ THz) with a Chern number of $|C|=1$ featuring a pair of edge modes traversing the phonon gap, as shown in Fig. 3(h). This provides definitive evidence for intrinsic phonon Chern states.
@@
Table 1: The direct phonon gaps at K points for the candidate materials, where $\gamma$ and $\Delta_{\text{z}}$ represent the phonon linewidth and the phonon Zeeman splitting, respectively. The observable phonon gap $\Delta_{\text{ph}}$ is defined as $\Delta_{\text{z}}-\gamma$.
@@
Discussion and conclusion. The intrinsic phonon Chern state, which hosts topologically protected edge phonon currents, holds significant potential for applications due to its operation without requiring strong external magnetic fields, Coriolis forces, or stress gradients [2, 49]. One potential advantage may be its ability to detect neutral particles by converting them into topological phonon edge currents. Neutral particles, especially dark matter particles, rarely participate in electromagnetic interactions, making their detection extremely challenging. However, if the energy transferred from the neutral particle to nucleus during the scattering or absorption falls within the first “conduction” band above the topological gap in the phonon dispersion, corresponding phonon modes can be excited [65, 66], driving the system out of equilibrium. As the system relaxes back to equilibrium, these high-frequency modes gradually shift to lower frequencies and necessarily cross the phonon Chern gap via topologically protected edge states, manifesting as a unidirectional edge phonon current immune to scattering [Fig. 4(a)]. This robust edge current is significantly more measurable than conventional delocalized bulk phonon excitations in materials. However, the operating temperature of detectors must be sufficiently low to prevent thermal excitation from overcoming the phonon gap, which would otherwise disrupt these edge currents. Thus, a large phonon Chern gap is highly desired to elevate the working temperature. Figure 4(b) lists phonon Chern gaps and transferred energy ranges of proposed candidate materials (see SM [44] for details), with transferred energy spanning approximately $20\sim 110\,\mathrm{meV}$. This energy range corresponds to low-energy dark matter which recently attracts much attention, therefore our work provides a potential detection mechanism for this category of dark matter candidates [65, 66].
@@
Figure 4: Detection of neutral particles based on the intrinsic phonon Chern state. (a) Schematic illustrations of the detection process. Left: Neutral particles (red beam) excite phonon modes (yellow circles in the lower panel) in high-frequency bands. Right: These phonon modes gradually relax to lower-frequency regimes via topologically protected edge phonon channels (orange and green lines across the gap), generating robust unidirectional phonon edge currents (orange and green arrows). (b) Transferred energy ranges and phonon Chern gaps from proposed candidate materials [Tab. S2 in SM [44]].
@@
In summary, we have established an ab initio theory enabling systematic predictions of EPC-induced magnetic properties of phonons and TRS-breaking phonon dispersions for insulators, semimetals and metals. Based on our theory, we construct the inertially decoupled model to seek intrinsic phonon Chern states in realistic materials, and propose candidates exhibiting EPC-induced large topological phonon gaps exceeding 0.3 THz. Furthermore, we demonstrate the potential capability of these topological phonon states to detect neutral particles through robust phonon edge currents. These advances could open broad pathways for investigating phonon magnetic effects in quantum materials.